OpenStudy (qqstory):

Complete the square. x^2+y^2-10x-23=0

OpenStudy (anonymous):

\[(x^2-10x+25)+y^2-23-25=0\]\[(x-5)^2+y^2=48\]\[(x-5)^2+y^2=(4\sqrt{3})^2\] Is this what you want?